Asia Europe North America
82 26 20
44 38 29
118 32 51
145 27 19
55 23 17
131 44
52 29
28

Find critical value, show me the steps, please.

Ata=0.10

Asia   Europe   North America  
82 26 20  
44 38 29  
118 32 51  
145 27 19  
55 23 17  
131 44      
52 29      
28      
Answer #1

Answer

We have to F critical value

we need degree of freedom numerator and denominator

df(numerator) = number groups -1

= 3-1............because we have three groups(asia, europe and north america)

=2

and

df(denominator) = total data values - number of groups

= (7+8+5) - 3

= 20-3

= 17

significance level = 0.10 (given)

Open F critical table and find values for alpha = 0.10

check df1 = 2 in the first row or top most row and check df2 = 17 in the left most column or first column

then select the intersecting cell

we get

F critical = 2.645 (3 decimals)